OpenStudy (anonymous):

Plz Help.. Will give medal and fan.. A bag contains 2 white marbles and 7 purple marbles. Two marbles are drawn at random. One marbles is drawn and not replaced, Then a second marbles is drawn. a). What is the probability of selecting a purple marble and then a white marble? b). What is the probability of selecting two white marbles? c). Is there a greater chance of selecting two white marbles in a row or two purple marbles in a row? Show your work..

OpenStudy (anonymous):

Hi again MathStudent55.. Could you help me?

OpenStudy (mathstudent55):

Ok. a) when you start, how many purple marbles are there and how many total marbles are there?

OpenStudy (anonymous):

so 7/9?

OpenStudy (mathstudent55):

Yes, for the first drawing of a marble. Now let's say you did draw a purple marble, and you don't replace the marble. For the second drawing in part a), how many purple marbles are there now, and how many total marbles are there now?

OpenStudy (anonymous):

I have this same question!

OpenStudy (mathstudent55):

@Madi_Lee Still there?

OpenStudy (anonymous):

I didn't get what you said MathStudent55

OpenStudy (mathstudent55):

In part a) we already drew one marble. The probability of the first marble being purple is 7/9 bec there were 7 purple marbles and 9 total marbles. Ok so far?

OpenStudy (anonymous):

Ok

OpenStudy (mathstudent55):

Now that we drew one purple marble from the bag, we need to draw a second marble, and we want it also to be purple.

OpenStudy (anonymous):

I think I got it now.. Thank you

OpenStudy (mathstudent55):

Since we took a purple marble out of the bag in the first drawing, how many purple marbles are left in the bag, and how many total marble are left in the bag for our second drawing?

OpenStudy (anonymous):

so 6/8

OpenStudy (mathstudent55):

Right. The probability of the two events happening is the product of the probabilities.

OpenStudy (mathstudent55):

P(two purple marbles) = 7/9 * 6/8

OpenStudy (mathstudent55):

That is part a)

OpenStudy (mathstudent55):

Of course, once you multiply the fractions, or even before, you can try to reduce them.

OpenStudy (anonymous):

so a). is 7/12?

OpenStudy (anonymous):

we multiply 7/9 x 6/8 = 7/12?

OpenStudy (mathstudent55):

Correct.

OpenStudy (anonymous):

ok so part a). the answer is 7/12

OpenStudy (mathstudent55):

For part b), do the same using the correct number of white marbles for the first and second drawings. The number of total marbles for both drawing will be the same as for the purple marbles.